OpenStudy (anonymous):

What is the formula for the sum of the first n positive even integers? What is it in Sigma notation?

OpenStudy (anonymous):



OpenStudy (anonymous):

I actually think the formula is n(n+1) by looking online. I just can't figure out how this is derived.

OpenStudy (anonymous):

oh, you are right! i now understand the question

OpenStudy (anonymous):

Do you know how that formula is derived?

OpenStudy (anonymous):

2+4+6+......+2n=2(1+2+3+....+n)=2*n(n+1)/2=n(n+1)  or